# Quantum Complexity of Nonlocal Field Theories

### v2

@@ 4.2 Action complexity (CA) for the isotropic case (5D)
    $S_{\mathcal{X}}$ $=\frac{-2}{16\pi G_{N}}\int d^{3}X\,\sqrt{\gamma}\ln\left|\frac{n_{+}\cdot n_{-}}{2}\right|\Bigg|_{r=r_{b}},$
    $=\frac{-2V_{3}r_{b}^{3}(1+a^{4}r_{b}^{4})^{1/2}}{16\pi G_{N}}\ln\left(\frac{\alpha^{2}\left(a^{4}r_{b}^{4}+1\right){}^{8/3}}{r_{b}^{2}}\right).$
@@
    $S_{\mathcal{X}}$ $=\frac{V_{3}r_{b}^{3}}{16\pi G_{N}}\left(4\ln\left(\frac{r_{b}}{\alpha}\right)+O(a^{4})\right).$ (118)
@@
    $S_{\mathcal{X}}$ $=\frac{V_{3}r_{b}^{3}}{16\pi G_{N}}\left(-\frac{64}{3}a^{2}r_{b}^{2}\ln\left(ar_{b}\right)+4a^{2}r_{b}^{2}\ln\left(\frac{r_{b}}{\alpha}\right)\right)+O(a^{-2}).$ (119)
@@
Summing over all the perturbative contributions obtained when the UV cutoff scale dominates the non-commuative coupling i.e. $a\ll 1/r_{b}$, consistently reproduces the total action complexity for the known case of pure AdS to be
$\mathcal{C}$ $=\frac{V_{3}r_{b}^{3}}{16\pi G_{N}}\left(4\ln 3+O(a^{4})\right)$ (120)
Here again, consistency with the known result Reynolds:2016rvl fixes $\alpha=l_{ct}$.
@@
On the other hand, in the extreme noncommutative regime i.e. $a\gg 1/r_{b}$ the action complexity is given by the sum of (112), (115), (117) and (119) to be
$\mathcal{C}$ $=\frac{a^{2}r_{b}^{5}V_{3}}{16\pi G_{N}}\left(-\frac{64}{3}\ln\left(a\right)-\frac{58}{3}\ln\left(r_{b}\right)+2\ln\left(5\right)+\frac{4}{3}\right)+O(1).$ (121)
Here we, as we did in the previous sections, in the end we have rescaled every argument of the logarithm in the units of AdS radius. Moreover, we have also set $l_{ct}$ to be one in the AdS radius scale without tempering the universal piece coming from the known pure AdS result. It is evident that the overall action complexity for the noncommutative case is negative unless the noncommutativity parameter satisfies the bound $\ln\left(\frac{a^{64}r_{b}^{58}}{5^{6}}\right)<4$ or $\left(\frac{a^{64}r_{b}^{58}}{5^{6}}\right)<e^{4}$. Moreover, the leading order scaling with the noncommutative coupling is the product of quadratic and logarithmic scaling with the nonlocal coupling parameter. It is however clear that due to the involvement of the UV scale the complexity cannot be made positive in any range of values in the coupling space. This is hinting at the theory being UV incomplete, i.e. it has UV instability (in past work Katoch:2023dfh we have also noted that negative holographic complexity could imply a system without a stable ground state). Also, the action complexity in the noncommutative case scales with the fifth power of the UV cutoff, unlike the local field theory that is expected to scale cubically in the UV cutoff scale. This power law departure from the extensive scaling is interpreted as the effect of nonlocality on the action complexity.
@@ 5 Discussions and Outlook
Now we discuss the subregion volume complexity results. For the DDSYM case, we clearly observe a phase transition, mirroring the behavior of holographic EE (as well as other probes such as Wilson loops and low order correlation functions as have been shown in the past literature). The critical subregion size for both subregion complexity and entanglement entropy (worked out in the appendix) turns out to be the same thereby establishing a unanimous phase transition scale (length) for the onset of nonlocality (dipole interaction dominated physics). For subregions larger than the critical subregion, the subregion volume complexity scales linearly with subregion extension along the dipole deformation (same as what is expected for a local SYM theory), however for subregions smaller than the critical size (length), the subregion volume complexity scales cubically with the subregion extension along the dipole deformation. For the anisotropic NCSYM theory, we also observe a phase transformation with similar traits. There are two phases: for subregions larger than the critical subregion, the subregion volume complexity scales linearly with subregion extension along the noncommutative deformation (same as what is expected for a local SYM theory), however for subregions smaller than the critical size (length), the subregion volume complexity scales cubically with the subregion extension along the noncommutative deformation. The phase transition point, i.e. the critical subregion length can be extracted from the plots which agrees within numerical errors with the analytical expression obtained from the entanglement entropy transition point. These results are very similar the LST counterparts we studied previously Chakraborty:2020fpt ; Katoch:2022hdf where the subregion volume complexity exhibited the Hagedorn-like phase transition set the scale of the deformation parameter(s). For the isotropic NCSYM we were unable to get numerical plots to work out reliably and conclusively as there were runaway errors even for very small values of the noncommutative deformation parameter $a$. This is true not just for the subregion volume complexity but also for the holographic entanglement entropy. We plan to revisit this issue in the future.
@@
Next we discuss the action complexity results. The action complexity results in the weakly nonlocal (near AdS) limit are in sync with the volume complexity, i.e. leading UV divergence piece match (upto numerical factors involving spacetime dimensions). The subleading pieces, however are not in one-to-one correspondence with the subleading divergences in volume complexity. This is not a surprise and has been documented in the literature since the initial days of investigations into holographic complexity. Now one generic feature of the action complexity results, inferred from by inspecting the far AdS/ highly nonlocal limit of the action complexity. For the action complexity to remain positive, the nonlocal deformation parameter(s) must have an upper bound. This is reminiscent of our past investigations of Little String theories (LST) Katoch:2022hdf where the nonlocal irrelevant deformation couplings had to satisfy bounds for the complexity to remain real positive. There the bound reflected the fact that the LST with the irrelevant deformations was not UV complete. We believe a similar scenario at play here: the nonlocal theories we consider might not be UV complete for arbitrary large values of the nonlocal deformation parameter. For the DDSYM this upper bound puts a restriction on the ’t-Hooft coupling! This is a rather unique phenomenon as we have pointed out earlier - we are strictly operating in the SUGRA limit i.e. infinite ‘t-Hooft coupling regime, and action complexity leads to an upper bound on the ‘t-Hooft coupling! There is no precedent of such a phenomenon in the AdS/CFT literature where a SUGRA calculation ends up imposing a bound on the ‘t-Hooft coupling. For the NCSYM theory, both the anisotropic and isotropic case, examining the action complexity in the extreme noncommutative regime and demanding the complexity remain positive leads to an upper bound on the noncommutativity parameter (to be precise, the dimensionless ratio of the noncommutativity parameter and AdS radius). However, there is a clear distinction between the anisotropic and isotropic case: for the anisotropic case the upper bound is independent of the UV cutoff, while for the isotropic case the bound is a function of the UV cutoff! In the other limit, i.e. weakly noncommutative regime, the action-complexity leads to perturbative corrections to the local SYM theory (pure AdS) action complexity. Although physically not that interesting, this furnishes a nice consistency check of our action complexity results (they all reproduce the local SYM theory results in the weak deformation regime). We also want to point out to a technical subtlety regarding the action complexity computations. For the action complexity computations, we have been using a 10d type IIB SUGRA action and in such case one has to incorporate a topological term Kurlyand:2022vzv ; Apolo:2025wcl . Otherwise the action complexity does not reproduce the expected answer in the pure AdS limit. If instead one uses a dimensionally reduced effective 5d SUGRA action, then one does not need to incorporate this term.
@@ Acknowledgements
We thank the organizers and participants of the BIRS-CMI workshop on “Quantum Gravity and Information Theory: Modern Developments” (9-14 Nov. 2025) which offered us a chance to present our results. We also thank Juan F. Pedraza for drawing our attention to the relevant works Fischler:2013gsa ; Couch:2017yil ; Eccles:2021zum . The work of SP and SR is supported by the CRG grant of the Anusandhan National Research Foundation (ANRF), Department of Science and Technology (DST), CRG/2023/001120 (“Many facets of complexity: From chaos to thermalization”). The work of SR is also supported by the IIT Hyderabad seed grant SG/IITH/F171/2016-17/SG-47 as well the IITH funds RDF/IITH/F171/SR. The work of GK is supported by IIT Indore Sponsored Research Project no. CRG/2023/000904 (“From Modular Hamiltonian to Operator Algebra in Gauge/Gravity Dictionary”).

### v3

@@ 4.2 Action complexity (CA) for the isotropic case (5D)
    $S_{\mathcal{X}}$ $=\frac{-2}{16\pi G_{N}}\int d^{3}X\,\sqrt{\gamma}\ln\left|\frac{n_{+}\cdot n_{-}}{2}\right|\Bigg|_{u=u_{b}},$
    $=\frac{-2V_{3}u_{b}^{3}(1+a^{4}u_{b}^{4})^{1/2}}{16\pi G_{N}}\ln\left(\frac{\alpha^{2}\left(a^{4}u_{b}^{4}+1\right){}^{8/3}}{u_{b}^{2}}\right).$
@@
    $S_{\mathcal{X}}$ $=\frac{V_{3}u_{b}^{3}}{16\pi G_{N}}\left(4\ln\left(\frac{u_{b}}{\alpha}\right)+O(a^{4})\right).$ (120)
@@
    $S_{\mathcal{X}}$ $=\frac{V_{3}u_{b}^{3}}{16\pi G_{N}}\left(-\frac{64}{3}a^{2}u_{b}^{2}\ln\left(au_{b}\right)+4a^{2}u_{b}^{2}\ln\left(\frac{u_{b}}{\alpha}\right)\right)+O(a^{-2}).$ (121)
@@
Summing over all the perturbative contributions obtained when the UV cutoff scale dominates the non-commuative coupling i.e. $a\ll 1/u_{b}$, consistently reproduces the total action complexity for the known case of pure AdS to be
$\mathcal{C}_{\mathcal{A}}$ $=\frac{V_{3}}{16\pi^{2}G_{N}\epsilon^{3}}\left(4\ln\left(\frac{3l_{ct}}{\alpha}\right)\right)+O(a^{4}).$ (122)
Where in the end, we have expressed $u_{b}=\frac{1}{\epsilon}.$
@@
On the other hand, in the extreme noncommutative regime i.e. $a\gg 1/u_{b}$ the action complexity is given by the sum of (114), (117), (119) and (121) to be
$\mathcal{C}_{\mathcal{A}}$ $=\frac{a^{2}V_{3}}{16\pi^{2}G_{N}\epsilon^{5}}\left(-\frac{64}{3}\ln\left(a\right)+\frac{58}{3}\ln\left(\epsilon\right)+2\ln\left(5\right)+\frac{4}{3}+\ln(l_{ct})-4\ln(\alpha)\right)+O(1).$ (123)
Here we, as we did in the previous sections, in the end we have rescaled every argument of the logarithm in the units of AdS radius. It is evident that the overall action complexity for the noncommutative case is negative unless the noncommutativity parameter satisfies the bound $\ln(\frac{a^{64}\alpha^{12}}{5^{6}l_{ct}^{3}\epsilon^{58}})<4$. Also, the leading order scaling with the noncommutative coupling is the product of quadratic and logarithmic scaling with the nonlocal coupling parameter. It is however clear that due to the involvement of the UV scale the complexity cannot be made positive in any range of values in the coupling space. This is hinting at the theory being UV incomplete, i.e. it has UV instability (in past work Katoch:2023dfh we have also noted that negative holographic complexity could imply a system without a stable ground state). Also, the action complexity in the noncommutative case scales with the fifth power of the UV cutoff, unlike the local field theory that is expected to scale cubically in the UV cutoff scale. This power law departure from the extensive scaling is interpreted as the effect of nonlocality on the action complexity.
@@ 5 Discussions and Outlook
Now we discuss the subregion volume complexity results. For the DDSYM case, we clearly observe a phase transition, mirroring the behavior of holographic EE (as well as other probes such as Wilson loops and low order correlation functions as have been shown in the past literature). The critical subregion size for both subregion complexity and entanglement entropy (worked out in the appendix) turns out to be the same thereby establishing a unanimous phase transition scale (length) for the onset of nonlocality (dipole interaction dominated physics). For subregions larger than the critical subregion, the subregion volume complexity scales linearly with subregion extension along the dipole deformation (same as what is expected for a local SYM theory), however for subregions smaller than the critical size (length), the subregion volume complexity scales cubically with the subregion extension along the dipole deformation. For the anisotropic NCSYM theory, we also observe a phase transformation with similar traits. There are two phases: for subregions larger than the critical subregion, the subregion volume complexity scales linearly with subregion extension along the noncommutative deformation (same as what is expected for a local SYM theory), however for subregions smaller than the critical size (length), the subregion volume complexity scales cubically with the subregion extension along the noncommutative deformation. The phase transition point, i.e. the critical subregion length can be extracted from the plots which agrees within numerical errors with the analytical expression obtained from the entanglement entropy transition point. These results are very similar the LST counterparts we studied previously Chakraborty:2020fpt ; Katoch:2022hdf where the subregion volume complexity exhibited the Hagedorn-like phase transition set the scale of the deformation parameter(s). For the isotropic NCSYM we were unable to get numerical plots to work out reliably and conclusively as there were runaway errors even for very small values of the noncommutative deformation parameter $a$ (these are discussed in details in appendix B). This is true not just for the subregion volume complexity but also for the holographic entanglement entropy. We plan to revisit this issue in the future.
@@
Next we discuss the action complexity results. The action complexity results in the weakly nonlocal (near AdS) limit are in sync with the volume complexity, i.e. leading UV divergence piece match (upto numerical factors involving spacetime dimensions). The subleading pieces, however are not in one-to-one correspondence with the subleading divergences in volume complexity. This is not a surprise and has been documented in the literature since the initial days of investigations into holographic complexity. Now one generic feature of the action complexity results, inferred from by inspecting the far AdS/ highly nonlocal limit of the action complexity. For the action complexity to remain positive, the nonlocal deformation parameter(s) must have an upper bound. This is reminiscent of our past investigations of Little String theories (LST) Katoch:2022hdf where the nonlocal irrelevant deformation couplings had to satisfy bounds for the complexity to remain real positive. There the bound reflected the fact that the LST with the irrelevant deformations was not UV complete. We believe a similar scenario at play here: the nonlocal theories we consider might not be UV complete for arbitrary large values of the nonlocal deformation parameter. For the DDSYM this upper bound puts a restriction on the ’t-Hooft coupling! For the NCSYM theory, both the anisotropic and isotropic case, examining the action complexity in the extreme noncommutative regime and demanding the complexity remain positive leads to an upper bound on the noncommutativity parameter (to be precise, the dimensionless ratio of the noncommutativity parameter and AdS radius). However, there is a clear distinction between the anisotropic and isotropic case: for the anisotropic case the upper bound is independent of the UV cutoff, while for the isotropic case the bound is a function of the UV cutoff! In the other limit, i.e. weakly noncommutative regime, the action-complexity leads to perturbative corrections to the local SYM theory (pure AdS) action complexity. Although physically not that interesting, this furnishes a nice consistency check of our action complexity results (they all reproduce the local SYM theory results in the weak deformation regime). We also want to point out to a technical subtlety regarding the action complexity computations. For the action complexity computations, we have been using a 10d type IIB SUGRA action and in such case one has to incorporate a topological term Kurlyand:2022vzv ; Apolo:2025wcl . Otherwise the action complexity does not reproduce the expected answer in the pure AdS limit. If instead one uses a dimensionally reduced effective 5d SUGRA action, then one does not need to incorporate this term.
Next we discuss some limitations of or exceptions to our particular approach and the broader framework. In the entirety of our work we have assumed that the leading UV divergence structure, in particular the scaling exponents, of complexity are unambiguous indicators of whether the underlying theory has local or nonlocal dynamics: volume extensive scaling indicating locality while hypervolume scaling indicating nonlocality. And indeed in the past literature, this approach has been utilized for holographic probes such as sugra correlators, Wilson loops, Entanglement entropy (extremal surfaces), this appears to the case, the scaling exponents change as one moves from a local theory in the IR to a nonlocal theory in the UV. However, when working with a probe such as complexity, one has to be extra cautious and bear in mind that are a multitude of different prescriptions for defining quantum complexity given a Hilbert space of a quantum system, and there is no guarantee the leading UV divergence will be identical for all of them. So while working complexity, it is entirely plausible that leading UV divergences are also distinguishing different prescriptions for complexity (both in the field-theory complexity), rather than distinguishing actual physical phases governed by local/nonlocal physics. This feature is unique to complexity, as other probes such as CFT correlators, or Wilson loops or entanglement entropy are defined unambiguously in the boundary quantum theory. Another, related issue in holographic setups is regarding the identification or interpretation of local dynamics in the gravity (bulk) with nonlocal quantum computation in the dual field theory May:2019yxi using shared entanglement. As such, the complexity of computational tasks in the boundary involving nonlocal quantum computation might not scale extensively in the boundary at all, refer to Bluhm:2025qgd ; Allerstorfer:2023ycc ; Asadi:2024fda for some work on the complexity theory for nonlocal quantum computations in quantum mechanics and field theory in the context of holography.
@@ Acknowledgements
We thank the organizers and participants of the BIRS-CMI workshop on “Quantum Gravity and Information Theory: Modern Developments” (9-14 Nov. 2025) which offered us a chance to present our results. We also thank Juan F. Pedraza for drawing our attention to the relevant works Fischler:2013gsa ; Couch:2017yil ; Eccles:2021zum . Finally, we thank the anonymous referee for several insightful comments which improved our manuscript appreciably and for drawing our attention to the highly interesting works Allerstorfer:2023ycc ; Asadi:2024fda ; Bluhm:2025qgd . The work of SP and SR is supported by the CRG grant of the Anusandhan National Research Foundation (ANRF), Department of Science and Technology (DST), CRG/2023/001120 (“Many facets of complexity: From chaos to thermalization”). The work of SR is also supported by the IIT Hyderabad seed grant SG/IITH/F171/2016-17/SG-47 as well the IITH funds RDF/IITH/F171/SR. The work of GK is supported by IIT Indore Sponsored Research Project no. CRG/2023/000904 (“From Modular Hamiltonian to Operator Algebra in Gauge/Gravity Dictionary”).
@@ Appendix A Entanglement entropy results for dipole deformed SYM
Appendix B Numerical errors in Isotropic NCSYM
If we proceed to study subregion complexity in the isotropic NCSYM case, the first step is to find the RT surface, which we take $\rho=\rho(z)$, so we will have the area as,
$A=4\pi l^{3}\int_{0}^{z_{0}}\frac{dz}{z^{3}}\Phi\rho^{2}\sqrt{\Phi^{2}+\rho^{\prime 2}}$ (128)
After performing variational analysis over this area functional we have the second order DE with the boundary conditions as $\rho(z_{0})=0$, and $\rho^{\prime}(z_{0})=\infty$ as,
$\rho^{\prime\prime}-\frac{2\left(1+\frac{a^{4}}{z^{4}}+\rho^{\prime 2}\right)}{\rho}-\frac{\left(5a^{4}z^{3}+3z^{7}\right)\rho^{\prime 3}}{(a^{4}+z^{4})^{2}}-\frac{3\rho^{\prime}}{z}=0$ (129)
Solving this equation numerically, with the boundary conditions we found that the solution is not compatible with the pure AdS limit. To resolve this issue we focus on the AdS case ($a\rightarrow 0$) for which we have the DE as,
$\rho^{\prime\prime}-\frac{2(1+\rho^{\prime 2})}{\rho}-\frac{3\rho^{\prime 3}}{z}-\frac{3\rho^{\prime}}{z}=0.$ (130)
It is known that the solution for this DE is $\rho^{2}+z^{2}=b^{2}$ Ryu:2006bv .
Upon numerically solving this equation, we find that the solution takes the form of a semicircle only within a certain range of the parameter $z_{0}$. Outside this range, the solution deviates from a semicircular profile, as illustrated in Fig. 4.
Figure 4: The extremal surface with extremal points $z_{0}=400,490,500$ respectively
This deviation is also reflected in the slope of the numerical plot of $\ln(A)$ versus $\ln(l)$. While the expected slope is 2, as the area $A\sim l^{2}$ for the pure AdS case. Our numerical results show a clear departure from this value, as demonstrated in the figure below.
Figure 5: The blue curve which has been plotted for the range of $z_{0}$ over which we get semi-circle solution we have slope two. For the orange curve which is for a broader range of $z_{0}$ we get slope value 3.